• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Holmes

    Log entry time 08:16:44 on June17,2004

    Entry number 126061

    keyword=Owl shift summary

    R2202 production 28 uA
     - Beam off 10 min.
    R2203 dithering 28 uA
    R2204 production 28 uA
    Feedback problems: R2205-8 are junk
    R2209 production 28.5 uA
    R2210 production 28.5 uA
    R2211 production 28.5 uA
     - PITA change during this run (wrong sign at first; A_Q got big)
     - BLUMI HV changed during this run
    R2212 dithering 28.5 uA
     - Afterward we noticed that in this run and in R2203, energy
       modulation was broken.  Found amplitude was reset to zero.
       During R2213 we changed amplitude to correct value of 100. 
    R2213 production 28.5 uA
    R2214 dithering 28.5 uA Energy modulation working.
    R2215 production 28.5 uA
     - Beam off 10 min.
    Total beam time in above runs = 403 minutes => 0.69 coulombs (0.35 effective 2-arm coulombs)
    R2216 production 28.5 uA started.